Find the total work done in moving a particle in a force field given by 𝐹 = 3π‘₯𝑦𝑖 βˆ’ 5𝑧𝑗 + 10π‘₯π‘˜ along the curve x = 𝑑 2 + 1, 𝑦 = 2𝑑2 , 𝑧 = 𝑑 3 π‘“π‘Ÿπ‘œπ‘š 𝑑 = 1 π‘‘π‘œ 𝑑 = 2.


1
Expert's answer
2021-05-28T05:40:57-0400

The total work done can be found as the integral of the scalar product between the force vector F and a certain displacement trajectory vector ds. We use the definition for the curve r as well to proceed:


"\vec{r} = \begin{pmatrix}
t^2+1, &2t^2, & t^3
\end{pmatrix}" (between t=1 and t=2) which gives"\frac{d\vec{r}}{dt} = \begin{pmatrix}
2t, &4t, & 3t^2
\end{pmatrix}"


"\vec{F} = \begin{pmatrix}
3xy, &-5z, & 10x
\end{pmatrix}" which we have to substitute in terms of the 't' variable:


"\vec{F} =\begin{pmatrix}
3(t^2+1)(2t^2), &-5(t^3), & 10(t^2+1)
\end{pmatrix}"


"\vec{F} = \begin{pmatrix}
6t^4+6t^2, &-5t^3, & 10+10t^2
\end{pmatrix}"


Then we substitute and calculate the total work:


"W=\int_{C}^{} \vec{F}\cdot d\vec{s} =\int_{C}^{} \vec{F}(\frac{\mathrm{d} \vec{r}}{\mathrm{d} t}) dt"


"W = \int_{1}^{2}

\begin{pmatrix}
6t^4+6t^2, &-5t^3, & 10+10t^2
\end{pmatrix}

\cdot

\begin{pmatrix}
2t, &4t, & 3t^2
\end{pmatrix} 
dt"


"=\int_{1}^{2} (12t^{5}+12t^{3}-20t^{4}+30t^{2}+30t^{4})dt"


"W=\int_{1}^{2} (12t^{5}+10t^{4}+12t^{3}+30t^{2})dt=\left [ 2t^{6}+2t^{5} +3t^{4}+10t^{3}\right ]_{1}^{2}= (2(2^6)+2(2^5)+3(2^4)+10(2^3))-(2(1^6)+2(1^5)+3(1^4)+10(1^3))=(128+64+48+80)-(2+2)+3+10)"


"W = (128+64+48+80)-(2+2+3+10) =320-17"


"W = 303\,u"


In conclusion, the total work done is 303 units.


Reference:

  • Young, H. D., Freedman, R. A., & Ford, A. L. (2006).Β Sears and Zemansky's university physicsΒ (Vol. 1). Pearson education.